The probabilities that a player will get 4-9 questions right on a trivia quiz are shown belovw PLEASE FIND MEAN, VARIANCE, AND STANDARD DEVIATION X 4 56 789 P(X) 0.06 0.2 0.4 0.1 0.14 0.1
Answer #1

Mean of the probability distribution is given as

\small E(X)=\sum xP(X=x)

Mean of the probability distribution that a player will get 4-9 questions right on a trivia quiz is

\small E(X)=4(0.06)+5(0.2)+6(0.4)+7(0.1)+8(0.14)+9(0.1)=6.36

Variance of a probability distribution is given as

\small Var(X)=E(X^{2})-\left ( E(X) \right )^{2}

\small E(X^{2})=\sum x^{2}P(X=x)

E(X2) 12(0.06) + 52(0.2) + 62(0.4) + 72(0.1) + 82(014) + 92(0.1) = 42.32

\small Var(X)=42.32-6.36^{2}=1.8704

Variance that a player will get 4-9 questions right on a trivia quiz is 1.87.

Standard deviation is the square root of the variance. So the standard deviation that a player will get 4-9 questions right on a trivia quiz is \small \sqrt{1.8704}=1.37

Mean = 6.36

Variance = 1.87

Standard deviation = 1.37